# On the Lavrentiev gap for manifold-valued maps

### v1

@@ 2.3. Musielak-Orlicz spaces
We conclude with the notion of vanishing web oscillations, which will be instrumental in the proof of Theorem 1.1. We recall that a set $F\subset\mathsf{M}$ is a web if it is compact, of Lebesgue measure zero, and its complement $\mathsf{M}\setminus F=\cup_{i=1}^{\bar{\textnormal{{m}}}}U_{i}$ consists of a finite number of components $U_{i}$ that are disjoint open connected sets.
@@
Let $\mathsf{M}$ be an oriented, compact Riemannian manifold and $w\in W^{1,\varphi}(\mathsf{M},\mathbb{R}^{N})$. We say that $w$ has vanishing web oscillations if for every $\varepsilon>0$ there exists a web $F\subset\mathsf{M}$ such that:
@@ 2.4. A few useful results
Let $\mathsf{M}$ be a Riemannian manifold as in (1.1). Let $\varphi(x,\cdot)$ be a Young function satisfying (1.3)-(1.7). Assume that either $\textnormal{{m}}=2$, or $\textnormal{{m}}\geq 3$ and
@@
Let $\mathsf{M}$ be a Riemannian manifold as in (1.1), with $\textnormal{{m}}\geq 2$. Let $\varphi(x,\cdot)$ be a Young function satisfying the $\Delta_{2}$-condition near infinity for every $x\in\mathsf{M}$. Assume that either (1.15) or (1.16) hold. Given any $\tilde{x}\in\mathsf{M}$, $R\in(0,R_{\mathsf{M}})$, any measurable function $w:\mathbb{B}_{R}^{\textnormal{{m}}}(\tilde{x})\to\mathbb{R}^{N}$ such that
@@
Let $\mathsf{M}$ be a Riemannian manifold as in (1.1), with $\textnormal{{m}}\geq 2$. Let $\varphi:\mathsf{M}\times[0,\infty)\to[0,\infty)$ be a function satisfying (1.3)-(1.8). Assume either (1.15) or (1.16) holds. Let $w\in(W^{1,\varphi}\cap L^{\infty})(\mathsf{M},\mathbb{R}^{N})$. Then, for every $\varepsilon>0$, there exists a finite family of geodesic spheres $\mathbb{S}^{\textnormal{{m}}-1}_{r_{i}}(x_{i})$, with radii $r_{i}<\varepsilon$, $i=1,\dots,i_{\varepsilon}$, and a continuous function $\eta:\mathsf{M}\to\mathbb{R}^{N}$ such that
@@ 3. Density results and absence of Lavrentiev phenomenon
$\int_{U_{h}}\frac{|\tilde{w}_{\varepsilon}-w|}{\text{diam}U_{h}}\,{\rm d}\mathscr{H}\leq c\int_{U_{h}}|D\tilde{w}_{\varepsilon}-Dw|\,{\rm d}\mathscr{H}\leq c\int_{U_{h}}|Dw|\,{\rm d}\mathscr{H},$
@@
(3.8) $\int_{\mathsf{M}}|\tilde{w}_{\varepsilon}-w|\,{\rm d}\mathscr{H}\leq c\,\varepsilon\,\lVert Dw\rVert_{L^{1}(\mathsf{M})},$
@@
whence $\lim_{\varepsilon\to 0}\|D\tilde{w}_{\varepsilon}\|_{L^{\varphi}(\mathsf{M})}=\|Dw\|_{L^{\varphi}(\mathsf{M})}$, which combined with (3.8), (3.9) and the uniform convexity of the space $L^{\varphi}$ (see for instance [36, Theorem 3.6.6.6]) yields
@@
$\int_{\mathsf{M}}\varphi(x,|D(\eta_{\varepsilon}(w))-Dw|)\,{\rm d}\mathscr{H}$ $=\int_{w^{-1}(O_{\varepsilon}T^{\textnormal{{k}}})}\varphi(x,|D(\eta_{\varepsilon}(w))-Dw|)\,{\rm d}\mathscr{H}\noindent$
$\leq c\left(\frac{1}{\varepsilon^{\gamma}}\int_{w^{-1}(O_{\varepsilon}T^{\textnormal{{k}}})}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}+\int_{w^{-1}(O_{\varepsilon}T^{\textnormal{{k}}})}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}\right)$
@@
(3.21) $\int_{\mathsf{M}}\varphi(x,|D(\eta_{\varepsilon}(w))-Dw|)\,{\rm d}\mathscr{H}\to 0.$
@@
(3.24) $\int_{\mathsf{M}}\varphi(x,|D(\pi\circ\tilde{w}_{{\ell}(\varepsilon)})-Dw|)\,{\rm d}\mathscr{H}<\varepsilon^{\gamma+1}.$
@@
$\varepsilon^{-\gamma}\int_{\mathsf{M}\setminus A_{\varepsilon}}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}$ $\leq\varepsilon^{-\gamma}\int_{w^{-1}(O_{2\varepsilon}T^{\textnormal{{k}}})}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}+\varepsilon^{-\gamma}\int_{v_{\varepsilon}^{-1}(\overline{O_{\varepsilon}T^{\textnormal{{k}}}})\setminus w^{-1}(O_{2\varepsilon}T^{\textnormal{{k}}})}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}$
$\qquad+\varepsilon^{-\gamma}\int_{v^{-1}_{\varepsilon}(\mathbb{R}^{N}\setminus\mathsf{N})}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}={\textnormal{{II}}_{1}}+{\textnormal{{II}}_{2}}+{\textnormal{{II}}_{3}}.$
@@
(3.31) $\int_{\mathsf{M}}\varphi(x,|D(\eta_{\varepsilon}(w))-Dw|)\,{\rm d}\mathscr{H}\leq c\lVert\varphi(\cdot,|Dw|)\rVert_{L^{1}(\mathsf{M})}.$
@@
(3.32) $\int_{\mathsf{M}}\varphi(x,|Dw_{\varepsilon}-D(\eta_{\varepsilon}(w))|)\,{\rm d}\mathscr{H}\leq\textnormal{{I}}+\textnormal{{II}}_{2}+\textnormal{{II}}_{3}+c\lVert\varphi(\cdot,|Dw|)\rVert_{L^{1}(\mathsf{M})}.$
@@ 4. Counterexample
This final section is devoted to proving the foundamental role played by assumption (1.8). When this condition fails, counterexamples arise. We consider the double phase functional (1.9). In this context condition (1.8) is implied by (1.10), where $\alpha$ is the Hölder exponent of the function $a(\cdot)$. Assuming (1.19) and taking the sphere $\textnormal{{S}}^{N-1}_{{\Lambda}}$ as target manifold, we construct a map $u\in u_{0}+W^{1,\varphi}((-1,1)^{\textnormal{{m}}},\textnormal{{S}}^{N-1}_{{\Lambda}})$ that cannot be approximated by smooth sphere-valued maps, where $\Lambda$ will be specified below. The construction, adapted from [2], provides the first vectorial counterexample and extends naturally to maps between manifolds. Note that, (1.10) is sharp when $p\leq\textnormal{{m}}$. We remark once again that the counterexample arises with a target manifold that is (N2)(N-2)-connected and even when $p\geq\textnormal{{m}}$, independently of the smoothness of the boundary data or the domain. Moreover, the functional has the Uhlenbeck structure, i.e. it is radial in the $z$-variable. So, take
@@
$\leq c\int_{0}^{1}t^{\frac{-\alpha+q(1-p_{0})}{q-1}}\mathcal{H}^{\textnormal{{m}}-1}\left(\,{\rm dist}(\cdot,\mathcal{C}^{\textnormal{{m}}-1}_{\lambda})\leq t/2\right)\,{\rm d}t$
@@
(4.28) $\mathrel{\makebox[24.33339pt][c]{$\displaystyle\overset{\mathrm{}}{=}$}}\sigma_{*}m_{*}\int_{Q_{\mathsf{M}}}\langle D\tilde{u}^{\textnormal{{S}}},\textnormal{{b}}\rangle\,{\rm d}x+\int_{Q_{\mathsf{M}}}\varphi(x,|Du_{0}^{\textnormal{{S}}}|)\,{\rm d}x-\frac{\sigma_{*}m_{*}}{2\sqrt{\frac{1}{4}+N-1}}-\tau$
(4.31) $\mathrel{\makebox[24.33339pt][c]{$\displaystyle\overset{\mathrm{}}{>}$}}\frac{3\sigma_{*}m_{*}}{8\sqrt{\frac{1}{4}+N-1}}+\int_{Q_{\mathsf{M}}}\varphi(x,|Du_{0}^{\textnormal{{S}}}|)\,{\rm d}x$
@@
(4.37) $\mathrel{\makebox[24.88895pt][c]{$\displaystyle\overset{\mathrm{}}{>}$}}\inf_{w\in W^{1,\varphi}_{\tilde{u}_{0}^{\textnormal{{S}}}}(Q_{\mathsf{M}},\textnormal{{S}}^{N-1}_{m_{*}})}\int_{Q_{\mathsf{M}}}\varphi(x,|Dw|)\,{\rm d}x;$
@@
$\int_{Q_{\mathsf{M}}}\langle D\tilde{u}^{\textnormal{{S}}},\textnormal{{b}}\rangle dx=\int_{\partial Q_{\mathsf{M}}}(\textnormal{{b}}\nu)\cdot u_{*}^{\textnormal{{S}}}\,\,{\rm d}S=\frac{1}{\sqrt{\frac{1}{4}+(N-1)}},$
@@
which contradicts (4.19). Taking $\Lambda=m_{*}$, $\tilde{u}^{\textnormal{{S}}}_{0}=u_{0}$ and $v=\bar{u}$, the proof in the case $p_{0}<\textnormal{{m}}$ is concluded. Let us now analyze the other two cases.
@@
Applying [2, Lemma 6] we get $\varphi^{*}(\cdot,|\textnormal{{b}}|)\in L^{1}(Q_{\mathsf{M}})$. Moreover, we observe that (4.14) still holds, i.e., $\int_{Q_{\mathsf{M}}}\varphi^{*}(\cdot,\sigma|\textnormal{{b}}|)\leq c_{1}\sigma^{q^{\prime}}<\infty$. The rest of the calculations follows as in the sub-dimensional case. This concludes the proof.
@@
(i) C.A. Antoninini was partly funded by GNAMPA of the Italian INdAM - National Institute of High Mathematics (grant number not available);

### v2

@@ 2.5. Vanishing web oscillations
Here we recall the notion of vanishing web oscillations, which will be instrumental in the proof of Theorem 1.1. We recall that a set $F\subset\mathsf{M}$ is a web if it is compact, of Lebesgue measure zero, and its complement $\mathsf{M}\setminus F=\cup_{i=1}^{\bar{\textnormal{{m}}}}U_{i}$ consists of a finite number of components $U_{i}$ that are disjoint open connected sets.
@@
Let $\mathsf{M}$ be an oriented, compact Riemannian manifold, $\partial\mathsf{M}=\emptyset$, and let $w\in W^{1,\varphi}(\mathsf{M},\mathbb{R}^{N})$. We say that $w$ has vanishing web oscillations if for every $\varepsilon>0$ there exists a web $F\subset\mathsf{M}$ such that:
@@ 2.6. A few useful results
Let $\mathsf{M}$ be a Riemannian manifold as in (1.1), with $\partial\mathsf{M}=\emptyset$. Let $\varphi(x,\cdot)$ be a Young function satisfying (1.3)-(1.7). Assume that either $\textnormal{{m}}=2$, or $\textnormal{{m}}\geq 3$ and
@@
Let $\mathsf{M}$ be a Riemannian manifold as in (1.1), with $\textnormal{{m}}\geq 2$ and $\partial\mathsf{M}=\emptyset$. Let $\varphi(x,\cdot)$ be a Young function satisfying the $\Delta_{2}$-condition near infinity for every $x\in\mathsf{M}$. Assume that either (1.15) or (1.16) hold. Given any $\tilde{x}\in\mathsf{M}$, $R\in(0,R_{\mathsf{M}})$, any measurable function $w:\mathbb{B}_{R}^{\textnormal{{m}}}(\tilde{x})\to\mathbb{R}^{N}$ such that
@@
Let $\mathsf{M}$ be a Riemannian manifold as in (1.1), with $\textnormal{{m}}\geq 2$ and $\partial\mathsf{M}=\emptyset$. Let $\varphi:\mathsf{M}\times[0,\infty)\to[0,\infty)$ be a function satisfying (1.3)-(1.8). Assume either (1.15) or (1.16) holds. Let $w\in(W^{1,\varphi}\cap L^{\infty})(\mathsf{M},\mathbb{R}^{N})$. Then, for every $\varepsilon>0$, there exists a finite family of geodesic spheres $\mathbb{S}^{\textnormal{{m}}-1}_{r_{i}}(x_{i})$, with radii $r_{i}<\varepsilon$, $i=1,\dots,i_{\varepsilon}$, and a continuous function $\eta:\mathsf{M}\to\mathbb{R}^{N}$ such that
@@ 3. Density results and absence of Lavrentiev phenomenon
Owing to Remark 2.1, it suffices to prove the theorem only in the case $\partial\mathsf{M}=\emptyset$.
@@
$\int_{U_{h}}\frac{|\tilde{w}_{\varepsilon}-w|}{\text{diam}U_{h}}\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\leq c\int_{U_{h}}|D\tilde{w}_{\varepsilon}-Dw|\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\leq c\int_{U_{h}}|Dw|\,{\rm d}\mathscr{H}^{\textnormal{{m}}},$
@@
(3.8) $\int_{\mathsf{M}}|\tilde{w}_{\varepsilon}-w|\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\leq c\,\varepsilon\,\lVert Dw\rVert_{L^{1}(\mathsf{M})},$
@@
whence $\lim_{\varepsilon\to 0}\|D\tilde{w}_{\varepsilon}\|_{L^{\varphi}(\mathsf{M})}=\|Dw\|_{L^{\varphi}(\mathsf{M})}$, which combined with (3.8), (3.9) and the uniform convexity of the space $L^{\varphi}$ (see for instance [36, Theorem 3.6.6]) yields
@@
In view of Remark 2.1, we just need to prove the theorem only in the case $\partial\mathsf{M}=\emptyset$.
@@
$\int_{\mathsf{M}}\varphi(x,|D(\eta_{\varepsilon}(w))-Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}$ $=\int_{w^{-1}(O_{\varepsilon}T^{\textnormal{{k}}})}\varphi(x,|D(\eta_{\varepsilon}(w))-Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\noindent$
$\leq c\left(\frac{1}{\varepsilon^{\gamma}}\int_{w^{-1}(O_{\varepsilon}T^{\textnormal{{k}}})}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}+\int_{w^{-1}(O_{\varepsilon}T^{\textnormal{{k}}})}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\right)$
@@
(3.21) $\int_{\mathsf{M}}\varphi(x,|D(\eta_{\varepsilon}(w))-Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\to 0.$
@@
(3.24) $\int_{\mathsf{M}}\varphi(x,|D(\pi\circ\tilde{w}_{{\ell}(\varepsilon)})-Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}<\varepsilon^{\gamma+1}.$
@@
$\varepsilon^{-\gamma}\int_{\mathsf{M}\setminus A_{\varepsilon}}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}$ $\leq\varepsilon^{-\gamma}\int_{w^{-1}(O_{2\varepsilon}T^{\textnormal{{k}}})}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}+\varepsilon^{-\gamma}\int_{v_{\varepsilon}^{-1}(\overline{O_{\varepsilon}T^{\textnormal{{k}}}})\setminus w^{-1}(O_{2\varepsilon}T^{\textnormal{{k}}})}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}$
$\qquad+\varepsilon^{-\gamma}\int_{v^{-1}_{\varepsilon}(\mathbb{R}^{N}\setminus\mathsf{N})}\varphi(x,|Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}={\textnormal{{II}}_{1}}+{\textnormal{{II}}_{2}}+{\textnormal{{II}}_{3}}.$
@@
(3.31) $\int_{\mathsf{M}}\varphi(x,|D(\eta_{\varepsilon}(w))-Dw|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\leq c\lVert\varphi(\cdot,|Dw|)\rVert_{L^{1}(\mathsf{M})}.$
@@
(3.32) $\int_{\mathsf{M}}\varphi(x,|Dw_{\varepsilon}-D(\eta_{\varepsilon}(w))|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\leq\textnormal{{I}}+\textnormal{{II}}_{2}+\textnormal{{II}}_{3}+c\lVert\varphi(\cdot,|Dw|)\rVert_{L^{1}(\mathsf{M})}.$
@@ 4. Counterexample
This final section is devoted to proving the fundamental role played by assumption (1.8). When this condition fails, counterexamples arise, and this is precisely the content of Theorem 1.4. We consider the double phase functional (1.9). In this context condition (1.8) is implied by (1.10), where $\alpha$ is the Hölder exponent of the function $a(\cdot)$. Assuming (1.19) and taking the sphere $\textnormal{{S}}^{N-1}_{{\Lambda}}$ as target manifold, we construct a map $u\in u_{0}+W^{1,\varphi}((-1,1)^{\textnormal{{m}}},\textnormal{{S}}^{N-1}_{{\Lambda}})$ that cannot be approximated by smooth sphere-valued maps, where $\Lambda$ will be specified below. The construction, adapted from [2], provides the first vectorial counterexample and extends naturally to maps between manifolds. Note that, (1.10) is sharp when $p\leq\textnormal{{m}}$. We remark once again that the counterexample arises with a target manifold that is (N2)(N-2)-connected and even when $p\geq\textnormal{{m}}$, independently of the smoothness of the boundary data or the domain. Moreover, the functional has the Uhlenbeck structure, i.e. it is radial in the $z$-variable. So, take
@@
$\leq c\int_{0}^{1}t^{\frac{-\alpha+q(1-p_{0})}{q-1}}\mathscr{H}^{\textnormal{{m}}-1}\left(\,{\rm dist}(\cdot,\mathcal{C}^{\textnormal{{m}}-1}_{\lambda})\leq t/2\right)\,{\rm d}t$
@@
(4.28) $\mathrel{\makebox[22.18062pt][c]{$\displaystyle\overset{\mathrm{}}{=}$}}\sigma_{*}m_{*}\int_{Q_{\mathsf{M}}}\langle D\tilde{u}^{\textnormal{{S}}},\textnormal{{b}}\rangle\,{\rm d}x+\int_{Q_{\mathsf{M}}}\varphi(x,|Du_{0}^{\textnormal{{S}}}|)\,{\rm d}x-\frac{\sigma_{*}m_{*}}{2\sqrt{\frac{1}{4}+N-1}}-\tau$
(4.31) $\mathrel{\makebox[22.18062pt][c]{$\displaystyle\overset{\mathrm{}}{>}$}}\frac{3\sigma_{*}m_{*}}{8\sqrt{\frac{1}{4}+N-1}}+\int_{Q_{\mathsf{M}}}\varphi(x,|Du_{0}^{\textnormal{{S}}}|)\,{\rm d}x$
@@
(4.37) $\mathrel{\makebox[22.61118pt][c]{$\displaystyle\overset{\mathrm{}}{>}$}}\inf_{w\in W^{1,\varphi}_{\tilde{u}_{0}^{\textnormal{{S}}}}(Q_{\mathsf{M}},\textnormal{{S}}^{N-1}_{m_{*}})}\int_{Q_{\mathsf{M}}}\varphi(x,|Dw|)\,{\rm d}x;$
@@
$\int_{Q_{\mathsf{M}}}\langle D\tilde{u}^{\textnormal{{S}}},\textnormal{{b}}\rangle\,{\rm d}x=\int_{\partial Q_{\mathsf{M}}}(\textnormal{{b}}\nu)\cdot u_{*}^{\textnormal{{S}}}\,\,{\rm d}S=\frac{1}{\sqrt{\frac{1}{4}+(N-1)}},$
@@
which contradicts (4.19). Taking $\Lambda=m_{*}$, $\tilde{u}^{\textnormal{{S}}}_{0}=u_{0}$ and $v=\bar{u}$, the proof of Theorem 1.4 in the case $p_{0}<\textnormal{{m}}$ is concluded. Let us now analyze the other two cases.
@@
Applying [2, Lemma 6] we get $\varphi^{*}(\cdot,|\textnormal{{b}}|)\in L^{1}(Q_{\mathsf{M}})$. Moreover, we observe that (4.14) still holds, i.e., $\int_{Q_{\mathsf{M}}}\varphi^{*}(\cdot,\sigma|\textnormal{{b}}|)\leq c_{1}\sigma^{q^{\prime}}<\infty$. The rest of the calculations follows as in the sub-dimensional case. This concludes the proof of Theorem 1.4.
@@
(i) C.A. Antonini was partly funded by GNAMPA of the Italian INdAM - National Institute of High Mathematics (grant number not available);